OpenStudy (anonymous):

I need help. -5over 3-4 over 3=?

OpenStudy (chris911):

what do u thini t is?

OpenStudy (richyw):

please type out the equation \[{-f \over 3-4} \over {3} \]\[\frac{-5}{3}-\frac{4}{3}\]

OpenStudy (richyw):

\[{-5 \over 3-4} \over {3}\]

OpenStudy (richyw):

\[\frac{-5}{\frac{3-4}{3}}\]

OpenStudy (anonymous):

\[- \frac{ 5 }{ 3 } -\frac{ 4 }{ 5 } =?\]

OpenStudy (richyw):

you want each fraction to have the same denominator. so multiply each one by one, but in a clever way.

OpenStudy (richyw):

\[-\frac{5}{3}\cdot 1-\frac{4}{5}\cdot 1\]\[=-\frac{5}{3}\cdot \frac{5}{5}-\frac{4}{5}\cdot\frac{3}{3}\]

OpenStudy (richyw):

\[=-\frac{25}{15}-\frac{12}{15}\]\[=-\frac{37}{15}\]

OpenStudy (richyw):

multiply each fraction* by one